OpenStudy (anonymous):

How many solutions are there to the following system of equations? 6x – 9y = –9 08_08_UT_07_a.gif A. infinitely many B. 2 C. 1 D. 0

OpenStudy (anonymous):

Please help I will give medals and fan!

OpenStudy (zpupster):

this is a line equation

OpenStudy (anonymous):

the bottom part of the equation is y=2/3x+1

OpenStudy (anonymous):

you can use substitution or elimination

OpenStudy (anonymous):

how do I do that @helpme1.2

OpenStudy (anonymous):

is it c.?

OpenStudy (zpupster):

no they are the same line you could solve it but if you notice that if you multiply y=2/3x+1 by 9 we get 9y =18/3x + 9

OpenStudy (anonymous):

So a.?

OpenStudy (zpupster):

yes A is correct